A boat attempts to travel straight across to the other side of a river. The boat can go 18 m/s and the river travels at 17 m/s. At what angle (to the nearest degree) does the boat need aim upstream to dock exactly on the other side of the river?
1
Expert's answer
2020-12-14T07:16:58-0500


Let "\mathbf{v}_w" be the vector of river velocity with magnitude "v_w = 17m\/s" and "\mathbf{v}" be the vector of boat velocity with magnitude "v = 18 m\/s". Than, adding these two vectors together as shown in the second figure above, obtain the vector "\mathbf{v}'" that should point North:


"\mathbf{v}' = \mathbf{v}_w + \mathbf{v}"

From the right triangle obtain angle "\theta" the boat needs aim upstream to dock exactly on the other side of the river:


"\sin\theta =\dfrac{v_w}{v} = \dfrac{17}{18}\\
\theta \approx 71\degree"

Answer. "71\degree".